Due to erratic sales of its sole product-a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financComplete this question by entering your answers in the tabs below. Req 1 Reg 2 Req3 Reg 4 Req 5A Req 5B Req 5C ...... The preComplete this question by entering your answers in the tabs below. Req 1 Reg 2 Req 3 Req 4 Req 5A Req 5B Req 5C Refer to theComplete this question by entering your answers in the tabs below. Req 1 Reg 2 Reg 3 Reg 4 Req 5A Req 5B Req 5C Refer to theComplete this question by entering your answers in the tabs below. Req 1 Reg 2 Req3 Req 4 Req 5A Req 5B Req 5C Refer to the oComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req3 Req 4 Req 5A Reg 5B Req 5C Refer to the oComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4 Req 5A Req 5B Req 5C Refer to theHelp solving
Answer #1

Solution 1:

CM ratio = Contribution margin / Sales = $100,800 / $252,000 = 40%

break-even point in unit sales = Fixed expenses / CM per unit = $112,800 / $8 = 14100 units

Break even point in dollar sales = Fixed expenses / CM ratio = $112,800 / 40% = $282,000

Solution 2:

increase (decrease) in the company’s monthly net operating income = Increase in contribution margin - Increase in advertising expenses

= ($88,000*40%) - $6,800 = $28,400

Solution 3:

New selling price per unit = $20*90% = $18

New CM per unit = $8 - $2 = $6 per unit

New fixed costs =$112,800 + $38,000 = $150,800

New sales volume = 12600*2 = 25200 units

New operating income = Contribution margin - Fixed costs = (25200*$6) - $150,800 = $400

Solution 4:

New contribution margin per unit = $8 - $0.70 = $7.30 per unit

Target income = $4,200

Nos of units to be sold to attain target profit = (Fixed cost + Net operating income) / CM per unit

= ($112,800 + $4,200) / $7.30 = 16027 units

Solution 5a:

New CM per unit = $8 +$3 = $11 per unit

New fixed costs = $112,800 + $52,000 = $164,800

New CM ration = $11 / $20 = 55%

New break even point in units = $164,800 / $11 = 14982 units

Breakeven point in dollar sales = $164,800 /55% = $299,636

Solution 5b:

Contribution format income statement - Operation Automated
Particulars Per unit % Total
Sales $20.00 100% $418,000.00
Variable costs $9.00 45% $188,100.00
Contribution margin $11.00 55% $229,900.00
Fixed costs $164,800.00
Net operating income $65,100.00
Contribution format income statement - Operation not Automated
Particulars Per unit % Total
Sales $20.00 100% $418,000.00
Variable costs $12.00 60% $250,800.00
Contribution margin $8.00 40% $167,200.00
Fixed costs $112,800.00
Net operating income $54,400.00

Solution 5c:

Yes, it is recommended company should automate its operations